### v1

@@ D.3   Dropout parameter selection
Figure 12 illustrates a comparison of the optimal values of penalty functions normalized by the estimated values obtained using the BS delta. It is noteworthy that the RNN-FNN network consistently outperforms the benchmark regardless of the dropout parameter value. Additionally, the selection of the dropout parameter remains consistent across all objective functions, with a dropout probability ($p$) of 50% yielding optimal performance. Consequently, for all our experiments, a dropout regularization parameter of $p=50\%$ is adopted. This parameter drives the likelihood of randomly dropping out a fraction of the units within a neural network during training, thereby generating varied architectures for each training iteration. As demonstrated in Warde-Farley et al., (2013), this method not only effectively mitigates overfitting but also boosts performance, which is consistent with our numerical results.
@@
  • Results are computed using 100,000 out-of-sample paths in the absence of transaction costs ($\kappa=0\%$). Agents are trained under the reduced state space $(\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R},\{h_{t,i}\}^{5}_{i=1})$ according to the conditions outlined in Section 4.3.1. All the metrics are expressed in proportion of values obtained under the Black-Scholes delta hedging strategy.
Online Appendix (not part of the paper)
D.4   JIVR Model parameters
Table 5: Estimated Gaussian copula parameters
$\epsilon_{t,R}$ 1.000
$\epsilon_{t,1}$0.550 1.000
$\epsilon_{t,2}$0.690 0.140 1.000
$\epsilon_{t,3}$ 0.030 -0.030 -0.0100 1.000
$\epsilon_{t,4}$0.220 0.250 0.120 0.280 1.000
$\epsilon_{t,5}$0.340 0.170 0.370 0.130 -0.050 1.000
Table 6: JIVR model parameter estimates
Parameter $\beta_{1}$ $\beta_{2}$ $\beta_{3}$ $\beta_{4}$ $\beta_{5}$ S&P500
$\theta_{1}$ 0.996290 -0.013869 0.002841
$\theta_{2}$ 0.003669 0.877813 0.001300
$\theta_{3}$0.032640 0.997071 0.003722 -0.004198
$\theta_{4}$ 0.980269
$\theta_{5}$0.047789 0.986019
$\nu$ 0.089445
$\sigma\sqrt{252}$ 0.380279 0.052198 0.048641 0.051536
$\omega$ 0.267589 0.977291
$\kappa$ 0.838220 0.965751 0.974251 0.945377 0.980844 0.888977
$a$ 0.134152 0.098272 0.092646 0.102201 0.100502 0.056087
$\gamma$0.111813 -1.482862 0.096766 0.060558 -0.102996 2.507796
$\zeta$ 0.143760 0.852943 0.029109 -0.159051 0.092664 -0.641306
$\varphi$ 1.351070 1.538928 2.284780 1.449977 1.428477 2.039669
D.5   Quadratic hedging problem
@@
  • These results are computed considering the hedging error of 99,000 out-of-sample independent paths from the Balck-Scholes market with yearly parameters $\mu=0.0892$ and and $\sigma=0.1952$. The RNN-FNN is trained based on 400,000 independent paths under the same scheme.

### v2

@@ E.3  Dropout parameter selection
  • Results are computed using 100,000 out-of-sample paths in the absence of transaction costs ($\kappa=0\%$). Agents are trained under the reduced state space $(\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R},\{h_{t,i}\}^{5}_{i=1})$ according to the conditions outlined in Section 4.3.1. All the metrics are expressed in proportion of values obtained under DH.
Figure 13 illustrates a comparison of the optimal values of penalty functions normalized by the estimated values obtained using DH. It is noteworthy that the RNN-FNN network consistently outperforms the benchmark regardless of the dropout parameter value. Additionally, the selection of the dropout parameter remains consistent across all objective functions, with a dropout probability ($p$) of 50% yielding optimal performance. Consequently, for all our experiments, a dropout regularization parameter of $p=50\%$ is adopted. This parameter drives the likelihood of randomly dropping out a fraction of the units within a neural network during training, thereby generating varied architectures for each training iteration. As demonstrated in Warde-Farley et al., (2013), this method not only effectively mitigates overfitting but also boosts performance, which is consistent with our numerical results.
Appendix F Quadratic hedging problem
@@
  • These results are computed considering the hedging error of 99,000 out-of-sample independent paths from the Black-Scholes market with yearly parameters $\mu=0.0892$ and and $\sigma=0.1952$. The RNN-FNN is trained based on 400,000 independent paths under the same scheme.
@@
Appendix G In-sample backtest
In this section, we benchmark our approach using a historical path of the JIVR model spanning from January 5, 1996, to December 31, 2020, to assess the effectiveness of RL agents. This experiment examines their performance based on the historical series $\{R_{t},\beta_{t}\}$. Specifically, we evaluate the hedging performance considering a new European ATM call option with a maturity of 63 days every 21 business days along this historical path. The option prices, serving as initial hedging portfolio values, are determined with the prevailing IV surface on the day the hedge is initiated.
  • Results are computed based on the observed P&L from hedging 296 simulated ATM Call options under real market conditions observed from May 1, 1996, to December 31, 2020. A new option is considered every 21 business days. Agents are trained under the reduced state space $(\delta_{t},\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R})$ according to the conditions outlined in Section 4.3.1.